OpenStudy (anonymous):

Find the vertex of the parabola whose equation is y = x2 - 4x + 6.

OpenStudy (anonymous):

pretty much the same as before, but this time \(-\frac{b}{2a}=-\frac{-4}{2}=2\)

OpenStudy (anonymous):

the 2 is an exponent. and thanks, these are the possible answers which one would it be? (-2, 18) (2, 2) (2, 6) @satellite73

OpenStudy (anonymous):

we know the first coordinate is 2 the second coordinate is \(2^2-4\times 2+6=4-8+6=-4+6=2\)

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

yw